List of papers produced by CLICS-II in year 3

This document lists all papers written by CLICS-II members, which were cited in Topic Progress Reports as deliverables.

[ABC95] M. Abadi, R. Bellucci, and P.-L. Curien. A model for formal parametric polymorphism: a per interpretation for system R, 1995. submitted for publication.

[AHS95] Thorsten Altenkirch, Martin Hofmann, and Thomas Streicher. Categorical reconstruction of a reduction free normalization proof. In David Pitt, David E. Rydeheard, and Peter Johnstone, editors, Springer LNCS 953, Category Theory and Computer Science, 6th International Conference, CTCS ’95, Cambridge, UK, August 1995.

[AM95a] S. Abramsky and G. McCusker. Games and full abstraction for the lazy λ-calculus. In D. Kozen, editor, Tenth Annual IEEE Symposium on Logic in Computer Science, pages 234–243. IEEE Computer Society Press, 1995.

[AM95b] S. Abramsky and G. McCusker. Games for recursive types. In C. Hankin, S. J. Gay, and R. Nagarajan, editors, Theory and Formal Methods ‘94. Imperial College Press, 1995.

[Acz95] Peter Aczel. Generalised set theory. In Proceedings of the 1994 Conference on Information-oriented appraoches to Logic, Language and Computation, CLSI Lecture Notes, 1995. Edited by Jerry Seligm and Dag Westerstahl.

[Aczar] Peter Aczel. Schematic consequence. In D.M. Gabbay, editor, What is a logical system? Oxford University Press, To appear.

[BBdP95] P. N. Benton, G. M. Bierman, and V. C. V. de Paiva. Computational types from a logical perspective I. Technical Report 365, Cambridge Univ. Computer Laboratory, May 1995.

[BD95] Ilya Beylin and Peter Dybjer. A short formal proof of coherence for monoidal categories. Draft paper, 1995.

[BR95] A. Bucalo and G. Rosolini. Repleteness and the associated sheaf, 1995. submitted for publication.

[Bie95] G. M. Bierman. What is a categorical model of intuitionistic linear logic? In 2nd Int. Conf. on Typed Lambda Calculus and Applications, Edinburgh, 1995, volume 902 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1995.

[Braa] T. Braüner. The Girard Translation Extended with Recursion. Full version of paper to appear in Proceedings of CSL ’94, LNCS 933, 1995. (Available by anon ftp from ftp.daimi.aau.dk at pub/BRICS/RS/95/13/BRICS-RS-95-13.ps.gz)

[CD] Thierry Coquand and Peter Dybjer. Intuitionistic model constructions and normalization proofs. Mathematical Structures in Computer Science. Accepted for publication. Preliminary version appears in Proceedings of the 1993 TYPES Workshop, Nijmegen.

[CN] A. Cheng and M. Nielsen. Open Maps (at) Work. April 1995. (Available by anon ftp from ftp.daimi.aau.dk at pub/BRICS/RS/95/23/BRICS-RS-95-23.ps.gz)

[CW] G-L. Cattani and G.Winskel. Presheaf models for concurrency. In preparation.

[Dyb95] Peter Dybjer. Internal type theory. In Proceedings of the Joint CLICS-TYPES Workshop on Categories and Type Theory, Göteborg, January 1995.

[EW95] Uffe H. Engberg and Glynn Winskel. Completeness results for linear logic on Petri nets. Accepted for Annals of Pure and Applied Logic, 1995.

[Gor95] Andrew D. Gordon. Bisimilarity as a theory of functional programming. In Eleventh Annual Conference on Mathematical Foundations of Programming Semantics, 1995.

[HM] J. M. E. Hyland and E. Moggi. The S-replete construction. In , editors, CTCS 1995, volume ? of Lecture Notes in Computer Science, pages ?, 1995.

[HO95] J. M. E. Hyland and C.-H. L. Ong. Pi-calculus, dialogue games and PCF. In Proceedings of 7th Annual ACM Conference on Functional Programming Languages and Computer Architecture, La Jolla, California. ACM, 1995.

[HO] J. M. E. Hyland and L. Ong. On full abstraction for PCF: I, II and III. preliminary draft.

[Hen95] B. Henhapl. Ein voll abstraktes, extensionales Modell für endliches PCF. Master’s thesis, Technische Hochschule Darmstadt, May 1995. 57pp.

[Hut95b] M. Huth. Knowledge organization in theories of mind. Knowledge Organization, 1, 1995. To appear.

[Hut95d] M. Huth. The Greatest Symmetric Monoidal Closed Category of Scottdomains. Technical Report 1734, Technische Hochschule Darmstadt, May 1995.

[JP95] A. Jung and H. Puhlmann. Types, logic, and semantics for nested databases. In M. Main and S. Brookes, editors, 11th Conference on Mathematical Foundations of Programming Semantics, volume 1 of Electronic Notes in Theoretical Computer Science. Elsevier Science Publishers B.V., 1995.

[Laf95] Y. Lafont, “Interaction Combinators”, submitted to publication, available by ftp anonymous on lmd.univ-mrs.fr in pub/lafont (file combinators.ps.Z)

[MS95] M. Marz and P. Selinger. A very well known theorem. Semsoc Newsletter, 4:7–8, 1995.

[Mar95] M. Marz. Monomorphe Typsysteme des λ-Kalküls. Master’s thesis, Technische Hochschule Darmstadt, June 1995. 46pp.

[Mog95] E. Moggi. Monadic semantics of CCS-like calculi. presented at the MASK workshop, Alghero, Italy, May 1995, 1995.

[Mogar] E. Moggi. Metalanguages and applications. In ??, volume ?? of CISM Lecture Notes. Springer Verlag, to appear. Revised and expanded version of lectures given at the Advanced School on Typed Lambda Calculus and Functional Programming (Udine, September 1994).

[NW] M.Nielsen and G.Winskel. Petri Nets and Bisimulations. January 1995. To appear in TCS. (ftp.daimi.aau.dk) (Available by anon ftp from ftp.daimi.aau.dk at pub/BRICS/RS/95/4/BRICS-RS-95-4.ps.gz)

[Oos94] Jaap van Oosten. Fibrations and Calculi of Fractions. November 1994. (Available by anon ftp from ftp.daimi.aau.dk at pub/BRICS/RS/94/37/BRICS-RS-94-37.ps.gz)

[Oos95] J. van Oosten. Basic Category Theory. January 1995.(Available by anon ftp from ftp.daimi.aau.dk at pub/BRICS/LS/95/1/BRICS-LS-95-1.ps.gz)

[Pit99] A. M. Pitts. Operationally-based theories of program equivalence. In P. Dybjer and A. M. Pitts, editors, Semantics and Logics of Computation. ? Cambridge University Press ?, 199? Based on lectures given at the CLICS-II Summer School on Semantics and Logics of Computation, Isaac Newton Institute for Mathematical Sciences, Cambridge UK, September 1995.

[Puh95] H. Puhlmann. Die Snack-Potenzkonstruktion: Grundlage einer Semantik genesteter Datenbanken. PhD thesis, Technische Hochschule Darmstadt, July 1995. 86pp.

[RP95] E. Ritter and A. M. Pitts. A fully abstract translation between a λcalculus with reference types and Standard ML. In 2nd Int. Conf. on Typed Lambda Calculus and Applications, Edinburgh, 1995, volume 902 of Lecture Notes in Computer Science, pages 397–413. Springer-Verlag, Berlin, 1995.

[Ros95] G. Rosolini. Notes on synthetic domain theory, 1995. available by ftp from theory.doc.ic.ac.uk in directory imported-papers/RosoliniG.

[SNW] V. Sassone, M. Nielsen, and G. Winskel. Models for Concurrency: Towards a Classification. To appear in Theoretical Computer Science.

[SR95] Th. Streicher and B. Reus. Continuation semantics : Control operators and abstract machines. Journal of Functional Programming, 1995. (Submitted.).

[Sas1] V. Sassone. An Axiomatization of the Algebra of Petri Net Concatenable Processes. To appear in Theoretical Computer Science.

[Sas2] V. Sassone. On the Category of Petri Net Computations. In Proceedings of the 6th International Conference on Theory and Practice of Software Development, TAPSOFT ‘95, LNCS n. 915, pp. 334–348, Springer-Verlag, 1995.

[Sima] H. Simmons. Axiomatic substitution. Preprint. Available on from the Dept of Computer Science, University of Manchester.

[Simb] H. Simmons. Boolean reflection of frames. Submitted, Available as Preprint.

[Sol1] S.Soloviev. Reductions in Intuitionistic Linear Logic Accepted for 4-th issue of Math.Structures in Comp. Sci., 1995.

[Sol] S.Soloviev. Proof of a S.Mac Lane conjecture. Submitted to ”Annals of Pure and Applied Logic”, extended abstract published in the Proceedings of CTCS-6, Cambridge, 1995.

[Sta95a] I. D. B. Stark. Names and higher-order functions. Technical Report 363, Cambridge Univ. Computer Laboratory, April 1995.

[Str95] C. Strüder. Stetige Operationen auf kontinuierlichen Bereichen. Master’s thesis, Technische Hochschule Darmstadt, April 1995. 38pp.

[Sün95a] Ph. Sünderhauf. Constructing a quasi-uniform function space. Topology and its Applications, 1995. To appear.

[Sün95b] Ph. Sünderhauf. A faithful computational model of the real numbers. Theoretical Computer Science, 1995. To appear.

[Sün95c] Ph. Sünderhauf. Quasi-uniform completeness in terms of Cauchy nets. Acta Mathematica Hungarica, 69(1–2), 1995. To appear.

[Tix95] R. Tix. Stetige Bewertungen auf topologischen Räumen. Master’s thesis, Technische Hochschule Darmstadt, June 1995. 51pp.

[Wag] K.R. Wagner. Liminf convergence in omega-categories. Submitted for publication.

[Win] G.Winskel. Hypercoherences. Manuscript, September 1994.